OpenStudy (anonymous):

When graphed in the (x,y) coordinate plane, at what point do the lines 2x + 3y = 5 and x = -2 intersect? A) (-2,0) B) (-2,5) C) (0,5/3) D) (0,5) E) (-2,3)

OpenStudy (anonymous):

confused help?

OpenStudy (juanpablojr):

have you tried google or looking it up?

OpenStudy (anonymous):

no...

OpenStudy (juanpablojr):

you might wanna check it out but ill help you anyways :)

OpenStudy (here_to_help15):

what number is this @PANTIEPOWER

OpenStudy (anonymous):

on my quiz its #4

OpenStudy (juanpablojr):

im stuck lol sorry

OpenStudy (anonymous):

its ok u tried tx tho

OpenStudy (texaschic101):

x = -2...this represents a vertical line....so the x value will always be -2....that eliminates a few answer choices so if x = -2...sub that into the other equation and find y 2x + 3y = 5 2(-2) + 3y = 5 -4 + 3y = 5 3y = 5 + 4 3y = 9 y = 3 your solution is : (2,3)

OpenStudy (anonymous):

u mean -2,3 :) so its E tx

OpenStudy (here_to_help15):

yup thats correct according to the link

OpenStudy (anonymous):

what link -.-

OpenStudy (here_to_help15):

http://www.analyzemath.com/practice_tests/act/act_sample_1.html this 1

OpenStudy (here_to_help15):

but they are not ordered the way urs might be u have to look around

OpenStudy (texaschic101):

oops..my bad...your right (-2,3)

OpenStudy (anonymous):

its fine thanks though :) and wow that link :0 im saving that